What is the energy of a photon that, when absorbed by a hydrogen atom, could cause an electronic transition from (a) the n=2 state to the n=5 state and (b) the n= 4 state to the n=6 state?
1
Expert's answer
2020-10-15T11:06:11-0400

Energy during level transition is given by :

a)

"\Delta E=13.6(\frac{1}{n_1^2}-\frac{1}{n_2^2})=13.6(\frac{1}{4}-\frac{1}{25})=13.6(\frac{21}{100})=2.85eV"


b)

"\Delta E=13.6(\frac{1}{n_1^2}-\frac{1}{n_2^2})=13.6(\frac{1}{16}-\frac{1}{36})=13.6(\frac{20}{576})=0.47eV"